OpenStudy (anonymous):

a bag contains 3 blue marbles, 9 green ,11 yellow marbles. Twice you draw a marble then replace it . Find p (blue then green) 27/529 27/23 15/529 12/23

OpenStudy (anonymous):

total marbles=3+9+11=23 \[required probability=\frac{ 3 }{23 }\times \frac{ 9 }{ 23}=\frac{ 27 }{529 }\]

OpenStudy (anonymous):

THANK YOU SO MUCH !

OpenStudy (anonymous):

@surjithayer